# Semi-orthogonality in Fukaya-Seidel mirrors to blowups of abelian varieties

### v1

@@ Abstract.
We prove a categorical homological mirror symmetry result for a blow-up of an abelian surface times $\mathbb{C}$, on the complex side. Specifically, the first author’s paper [Can20] provides evidence for homological mirror symmetry for genus 2 curves $\Sigma_{2}$ on the complex side, where $\Sigma_{2}$ is a hypersurface in an abelian surface $V\cong T^{4}$. To obtain a mirror to $\Sigma_{2}$, the generalized SYZ approach [HV00, AAK16] was used. That is, $(\text{Bl}_{\Sigma_{2}\times\{0\}}V\times\mathbb{C}_{y},y)=:(X,y)$ admits a Lagrangian torus fibration which has an SYZ mirror Landau-Ginzburg model $(Y,v_{0})$, known as the generalized SYZ mirror of $\Sigma_{2}$. The mirror to $X$ - without the superpotential - is obtained by removing a generic fiber from $(Y,v_{0})$. In this paper, we prove a categorical homological mirror symmetry result for $X$. To do so, we equip the symplectic mirror with a Fukaya category which involves both partial and full wrapping in the base of a symplectic Landau-Ginzburg model. Semi-orthogonality appears in the categorical invariants on both sides of HMS.
@@ 1.1. Background and motivation
Homological mirror symmetry (HMS) was conjectured by Kontsevich in [Kon95]. It is an equivalence, for certain pairs of complex manifold $X$ and symplectic manifold $Y$, of the bounded derived category of coherent sheaves and the Fukaya category [FOOO09], [Aur14], that is,
@@
The symplectic side is referred to as the “A-side” and the complex side as the “B-side”. So here $X$ is on the B-side and $Y$ is on the A-side. With the structure of a special Lagrangian torus fibration on a symplectic manifold, Strominger-Yau-Zaslow [SYZ96] formulated SYZ mirror symmetry, a geometric way to construct the complex mirror.
To find a mirror to a hypersurface $H$, the generalized SYZ construction can be used [HV00]. This is because a hypersurface doesn’t have an evident Lagrangian torus fibration (the way a toric or abelian variety $V$ does), so we instead consider the intermediary manifold $X$ given by the deformation to its normal cone, because that one does admit such a fibration. In this paper, we prove a categorical HMS result for the mirror $Y$ to the $X$ arising from the hypersurface $H$ given by a 1-parameter family of genus curves in a 4-torus $V$ (principally polarized abelian surface) considered in [Can20].
This overview of the generalized SYZ construction is based on [AAK16]. Let $V$ be a toric variety or a principally polarized abelian variety of complex dimension $n$. Let $\mathcal{L}\to V$ be an ample holomorphic line bundle with section $s$, and $H=s^{-1}(0)$ its divisor, a complex hypersurface of $V$. Let $X:=\text{Bl}_{H\times\{0\}}V\times\mathbb{C}$.
@@
$X$ is known as the deformation to the normal cone because the blow-up map composed with the projection to the $\mathbb{C}$ factor $y:X\to\mathbb{C}$ parametrizes a deformation of a smooth generic fiber to a singular fiber containing the exceptional divisor $E$ over $0\in\mathbb{C}$. Recall that the exceptional divisor is the projectivization of the normal bundle to the submanifold being blown-up, here $\mathbb{P}(\mathcal{N}_{H\times\{0\}\backslash V\times\mathbb{C}})$ where $\mathcal{N}_{H\times\{0\}\backslash V\times\mathbb{C}}\cong(\mathcal{L}\oplus\mathcal{O}_{\mathbb{C}})|_{H\times\{0\}}\to H\times\{0\}$.
@@
(1.2) $X=\text{Bl}_{H\times\{0\}}V\times\mathbb{C}=\{(x_{1},\ldots,x_{n},y,(u_{1}:u_{2}))\in\mathbb{P}(\mathcal{L}\oplus\mathcal{O})\mid s(x_{1},\ldots,x_{n})u_{2}=u_{1}y\}$
@@
We can canonically identify the preimage $y^{-1}(t)$ for $0\neq t\in\mathbb{C}$ with $V$ because for ${x}:=(x_{1},\ldots,x_{n})\in V$, then $u_{1}/u_{2}=s({x})/t$ is fixed. As outlined in Remark 1.1, over 0 we have a deformation to $y^{-1}(0)\cong V\cup_{H}E$ intersecting along $H$, as follows, see [AAK16, Proposition 7.3] for more details. The exceptional divisor $E$ is a $\mathbb{P}^{1}$-bundle
$\text{pr}_{V}\circ p:E\to H$
where $p:=\pi|_{E}$, given by
$E=\pi^{-1}(H\times\{0\})\ni({x},s({x})=0,(u_{1}:u_{2}))\mapsto{x}\in H$
for any $(u_{1}:u_{2})\in\mathbb{P}^{1}\cong\mathbb{P}(\mathcal{L}_{{x}}\oplus\mathcal{O}_{\{0\}})$. Also in $y^{-1}(0)$ we have the proper transform $\tilde{V}$ of $V\times\{0\}$, which is
$\tilde{V}=\overline{\pi^{-1}((V^{0}\backslash H)\times\{0\})}=\{({x},0,(1:0))\mid{x}\in V\}\cong V$
where $V^{0}$ is obtained by removing the toric divisors if any (not applicable for an abelian variety). Intersecting these two sets
@@
Remark 1.2 (Lagrangian torus fibration on $X$ [AAK16, §4]).
The moment map on $V\times\mathbb{C}$ is obtained from the torus action $(e^{i\theta_{1}}x_{1},\ldots,e^{i\theta_{n}}x_{n},e^{i\theta_{n+1}}y)$. Then using cut and paste operations as in [Sym03] on the image of the moment map, the moment polytope, this corresponds to blowing up $V\times\mathbb{C}$ along $H\times\{0\}$. The result is a new polytope that is the base of a Lagrangian torus fibration on $X$.
Generalized SYZ mirror symmetry. Here is a rough outline of the generalized SYZ procedure of [AAK16], which we will later describe in more detail for the setting of this paper. Let
$\Omega_{V\times\mathbb{C}}=i^{n+1}\bigwedge_{j=1}^{n}d\log x_{j}\wedge d\log y$
be a meromorphic $(n+1)$-form on $V\times\mathbb{C}$ with simple poles on the toric divisors $D_{V\times\mathbb{C}}=(V\times 0)\cup(D_{V}\times\mathbb{C})$ where $D_{V}=V\backslash V^{0}$, which is empty if $V$ is an abelian variety. Then we have a canonical form on $X$ given by
$\Omega=p^{*}\Omega_{V\times\mathbb{C}}$
@@
$D=\tilde{V}\cup p^{-1}(D_{V}\times\mathbb{C}).$
@@
For an abelian variety there are additional steps to quotient by the defining lattice. Both $X$ and $Y$ admit Kähler structures and the holomorphic functions $y:X\to\mathbb{C}$, $W_{0}:Y^{0}\to\mathbb{C}$, and $v_{0}:Y\to\mathbb{C}$ are Landau-Ginzburg models. In this paper, we take $X$ without superpotential to be the complex side (B-side) of mirror symmetry and $(Y^{0},W_{0})$ to be a symplectic Landau-Ginzburg model as defined in [ACLL23b, Definition 1.1].
@@
Related work. Evidence for categorical HMS of a 1-parameter family of complex genus 2 curves (as a hypersurface in the Jacobian torus) and their generalized SYZ mirror, was proved in [Can20] for a subcategory of the Fukaya category with fiber Lagrangians $\hat{\ell}_{k,[0]}$ and $\omega_{B}=0$ in the notation of [ACLL, §1.2.3 and Equation (2.36)]. SYZ mirror symmetry was proved in [KL19]. HMS for complex hypersurfaces of $(\mathbb{C}^{*})^{n}$ was proved in [AA24] and SYZ mirror symmetry in both directions for hypersurfaces of toric varieties in [AAK16, CLL12]. Another related work includes [Aur18], relating the Fukaya category of LG models of the form $((\mathbb{C}^{*})^{n},f)$ to the wrapped Fukaya category of a smooth fiber and of its complement in $(\mathbb{C}^{*})^{n}$; note that the hypersurface is on the A-side while here it is on the B-side. HMS for toric blow-ups has been studied in [HH22, §6.3], blowing up at toric fixed points, in contrast to hypersurfaces away from toric divisors as here. In the works [VWX23, Ker08] semi-orthogonal decompositions show up in the toric setting, as they do in our abelian variety setting here. Furthermore, [GKR17] discuss other versions of mirror symmetry for general type varieties.
@@ 1.2. Statement of main theorem
Generalized SYZ mirror symmetry for $V$ an abelian surface. We describe $X$ and $Y$ in more detail when $V$ is an abelian surface ($n=2$) and $H$ is a genus 2 curve. This is the setting of [Can20]; the HMS result there was for $H$ and $(Y,-v_{0})$. The HMS result here is for $X$ and $(Y^{0},W_{0})$.
Let $V=(\mathbb{C}^{*})^{2}/\tau\mathbb{Z}^{2}$ be the principally polarized abelian surface (topologically a 4-torus $T^{4}$) corresponding to the complex modular parameter $\tau=\frac{i}{2\pi}\log t\left(\begin{matrix}2&1\\ 1&2\end{matrix}\right)$, $0<t\in\mathbb{R}_{\ll 1}$, defined by quotienting by
$(\tau n)\cdot(x_{1},x_{2})=(e^{2\pi i\sum_{j=1}^{2}\tau_{1j}n_{j}}x_{1},e^{2\pi i\sum_{j=1}^{2}\tau_{2j}n_{j}}x_{2}),\quad(x_{1},x_{2})\in(\mathbb{C}^{*})^{2}.$
@@
Define $X:=\text{Bl}_{H\times\{0\}}V\times\mathbb{C}$ with blow-up map $\pi:X\to V\times\mathbb{C}$, exceptional divisor $p:E\to H\times\{0\}$, and inclusion $i:H\times\{0\}\to V\times\mathbb{C}$. To obtain the generalized SYZ mirror, we have an additional step of quotienting the above set-up by $\tau\mathbb{Z}^{2}$ after restricting $Y$ to a neighborhood where the group acts freely [Can20, Eqs. (3.10), (3.24), (3.25)]. For small $\varepsilon<1$,
@@
$\Delta_{Y}$ $=(\Delta_{\tilde{Y}})|_{\chi^{(0,0,1)}\leq\varepsilon}/\tau\mathbb{Z}^{2}$
where $\chi^{(0,0,1)}$ is the complex toric coordinate corresponding to the $\eta$ direction, $(0,0,1)$. We will denote it by $v_{0}$ following [Can20, AAK16]. The polytope $\Delta_{\tilde{Y}}$ looks like a bowl tiled by infinitely many hexagons and filled by an amount of $\varepsilon$. Recall from toric geometry that each vertex $v$ corresponds to a chart $\text{Spec}(\mathbb{C}[y_{1}^{v},y_{2}^{v},y_{3}^{v}])\cong\mathbb{C}^{3}$ in local inhomogeneous coordinates $y_{1}^{v},y_{2}^{v},y_{3}^{v}$, and edges define transition functions. The reason this polytope defines the SYZ mirror is because its vertices and edges match the SYZ charts and transition functions. Let
$v_{0}:=y_{1}^{v}y_{2}^{v}y_{3}^{v}.$
Then $(Y,-v_{0})$ is the generalized SYZ mirror to $H$. The relation between $(Y,v_{0})$ and $(Y^{0},W_{0})$ is
(1.4) $W_{0}=T^{\epsilon}(v_{0}-1)$
and
@@
where $T$ is the Novikov parameter and $\epsilon>0$ is small (different from the $\varepsilon$ in the domain of definition in Equation (1.3)). The Novikov parameter parametrizes the complex structure on $Y$ mirroring symplectic structures on $X$, the way $\tau$ parametrized in the other direction, the complex structure on $X$ mirroring symplectic structures on $Y$. Here $W_{0}:Y^{0}\to\mathbb{C}^{*}$ is a $T^{4}$-fibration over $\mathbb{C}^{*}$ with one singular fiber over $T^{\epsilon}$ (where $v_{0}=0$). We removed a smooth fiber from $Y$, so the singular fiber remains in $Y^{0}$. We use the symplectic form on $Y$ constructed in [Can20, §3.5], restricted to $Y^{0}$. The total space of $Y$ is smooth, and $\omega$ is modeled on the standard form for $\mathbb{C}^{3}$ near critical points of $v_{0}$ (the north and south pole of the banana manifold of three $\mathbb{P}^{1}$’s identified at their poles), while it is modeled on the standard product form for $\mathbb{CP}^{2}(3)\times\mathbb{D}$ away from the critical points.
@@
What was called $x,y,z$ in [Can20, Figure 6, §3.3] is $y_{1}^{(-1,-1,0)},y_{2}^{(-1,-1,0)},y_{3}^{(-1,-1,0)}$ here. And $\tau$ in [Can20] has been renamed $t$ here, in order to reserve $\tau$ for the more standard notation of the complex modular parameter.
To recap, here are the three forms of SYZ mirror symmetry in [AAK16] in the case of $V$ an abelian surface.
Theorem 1.5 ([AAK16, Theorem 10.4]).
@@
HMS for the third one is examined in [Can20]. HMS for the second one is the subject of this paper.
@@
In the case of an abelian variety (topologically a compact torus), note that since $W_{0}$ is a symplectic fibration with fibers given by closed, compact tori with volume, we cannot be in the exact setting by Stokes’ theorem. For symplectic form $\omega$, if $\omega|_{W_{0}^{-1}(z)}]=0$ then $\int_{W_{0}^{-1}(z)}\omega^{n}=0$ where $W_{0}^{-1}(z)\cong T^{2n}$, contradiction.
@@
Semi-orthogonal decomposition for blowups. We adapt the diagram at the start of [Orl92, §4] where their $\tilde{X}\to X$ is our blow up $\pi:\text{Bl}_{H\times\{0\}}V\times\mathbb{C}\to V\times\mathbb{C}$, their $p:\tilde{Y}\to Y$ is the exceptional divisor which is our $\mathbb{P}^{1}$-bundle $p:E\to H\times\{0\}$, and $j:E\to X$ is inclusion. That is, the following commutes:
@@
Then [Orl92, Assertion 4.2(b), Theorem 4.3] assert that there is a semi-orthogonal decomposition
@@
where $\mathcal{O}_{p}(-1):=\mathcal{O}_{E}(E)$. In the first factor, this means that a sheaf on $H$ is pulled back to the exceptional divisor and then extended by 0 to all of $X$.
This result holds, roughly speaking, because of Beilinson’s spectral sequence in [Huy06, Proposition 8.28] and the semi-orthogonality of $\mathcal{O}_{\mathbb{P}^{n}}$ with $\mathcal{O}(1)_{\mathbb{P}^{n}}$, namely that $H^{0}(\mathbb{P}^{n},\mathcal{O}(1))\cong\mathbb{C}[x_{0},\ldots,x_{n}]_{1}$ (degree 1 homogeneous polynomials in $(n+1)$-variables) but $H^{0}(\mathbb{P}^{n},\mathcal{O}(-1))=0$.
Main Theorem.
Definition 1.7 (notation $D^{b}_{\mathcal{L}}$, [Can20]).
Let $\text{pr}_{V}:V\times\mathbb{C}\to V$ and $\text{pr}_{\mathbb{C}}:V\times\mathbb{C}\to\mathbb{C}$ be the projections to each factor. We define the following subcategories.
@@
    $D^{b}_{\mathcal{L}}\mathrm{Coh}(X)$: full subcategory of objects in $j_{*}(p^{*}D^{b}_{\mathcal{L}}\mathrm{Coh}(H\times\{0\})\otimes\mathcal{O}_{E}(E))$ and $\pi^{*}D^{b}_{\mathcal{L}}\mathrm{Coh}(V\times\mathbb{C})$, and including shifts $[k]$

### v2

@@ Abstract.
We prove evidence of Kontsevich’s homological mirror symmetry conjecture (HMS) for a blow-up of an abelian surface times the complex plane, on the complex side, and its symplectic Landau-Ginzburg mirror. Specifically, the first author proved evidence of HMS for a 1-parameter family of genus 2 curves on the complex side, as a hypersurface in an abelian surface. The generalized SYZ mirror to the hypersurface is then the SYZ mirror to the Landau-Ginzburg model given by the blow-up of the abelian surface times the complex plane, along the hypersurface times zero, with superpotential given by projection to the complex plane. The mirror to the blow-up - without the superpotential - is obtained by removing a generic smooth fiber from the generalized SYZ mirror superpotential. We prove a categorical HMS result for the latter pair, between categories expected to split-generate. To do so, we equip the punctured superpotential with a Fukaya category which involves both partial and full wrapping in the base of the symplectic Landau-Ginzburg model due to the removal of a generic fiber. Semi-orthogonality appears in the categorical invariants on both sides of HMS.
Key words and phrases:
Homological mirror symmetry, Fukaya-Seidel category, symplectic Landau-Ginzburg model, derived category of coherent sheaves, semi-orthogonal decomposition, abelian varieties, generalized SYZ mirror symmetry, blow-up.
2020 Mathematics Subject Classification:
53D37 (primary), 18G70 (secondary)
C. Cannizzo was partially supported by NSF DMS-2316538. She also thanks the Collaborative Research Centre TRR 326 Geometry and Arithmetic of Uniformized Structures (GAUS) at Goethe Universität-Frankfurt for hosting her during the completion of this work.
@@ 1.1. Background and motivation
Homological mirror symmetry (HMS) was conjectured by Kontsevich [hms] as an equivalence of the Fukaya category [fooo1, fuk_intro] of a symplectic manifold $Y$ and the bounded derived category of coherent sheaves of a mirror complex manifold $X$,
@@
The symplectic side is referred to as the “A-side” and the complex side as the “B-side,” so here $X$ is on the B-side and $Y$ is on the A-side. One geometric way to construct the mirror is Strominger-Yau-Zaslow [syz] mirror symmetry. If one can find a special Lagrangian torus fibration on the symplectic manifold, the SYZ mirror is the T-dual torus fibration, which has the same base and dual fibers. Thus, SYZ mirror symmetry provides a candidate pair to prove HMS for. In this paper, we prove evidence in a new example that HMS holds for an SYZ mirror pair.
Calabi-Yau and Fano varieties, and their mirrors, were the first examples of HMS [fuk_abel, ab_sm, zp, sherid_CY, sherid_Fano, seidel, ueda_del_pezzo, Ab06, Ab09, FLTZ12]. Less is known about general type varieties and their mirrors. In particular, hypersurfaces of abelian and toric varieties have been studied more recently. To find a mirror to a hypersurface $H$, the generalized SYZ construction can be used [HV00, AAK]. This is because a hypersurface doesn’t have an evident Lagrangian torus fibration, the way a toric or abelian variety $V$ does. So we instead consider the intermediary manifold $X$ given by a deformation to the normal cone, because that does admit such a fibration. The superpotential, also known as a Landau-Ginzburg (LG) model, given by projection to the deformation parameter, has $H$ as the critical locus so the SYZ mirror superpotential is generalized SYZ mirror to $H$ [AAK, Ca20]. In this paper, we prove a categorical HMS result for $X$ without the superpotential, and its mirror.
Related work. Evidence for categorical HMS of a 1-parameter family of complex genus 2 curves as a hypersurface in the Jacobian torus was proved in [Ca20] for a subcategory of the Fukaya category. SYZ mirror symmetry was proved in [KL]. HMS for complex hypersurfaces of $(\mathbb{C}^{*})^{n}$ was proved in [AA24] and SYZ mirror symmetry for hypersurfaces of toric varieties in [AAK, CLL]. Other related work includes [Aur18], relating the Fukaya category of Landau-Ginzburg models $((\mathbb{C}^{*})^{n},f)$ to the wrapped Fukaya category of a smooth fiber and of its complement in $(\mathbb{C}^{*})^{n}$; note that the hypersurface is on the A-side while here it is on the B-side. HMS for toric blow-ups has been studied in [HH21, §6.3], blowing up at toric fixed points, in contrast to hypersurfaces away from toric divisors as here. The works [VWX23, Kerr08] consider semi-orthogonal decompositions in the toric setting, as we do in the abelian variety setting here. Furthermore, [GKR17] discuss other versions of mirror symmetry for general type varieties.
This overview of the generalized SYZ construction is based on [AAK, §3]. Let $V$ be a toric variety or a principally polarized abelian variety of complex dimension $n$. Let $\mathcal{L}\to V$ be an ample holomorphic line bundle with section $s$, and $H=s^{-1}(0)$ its divisor, a complex hypersurface of $V$. Let $X:=\text{Bl}_{H\times\{0\}}V\times\mathbb{C}$.
@@
$X$ is known as the deformation to the normal cone because the blow-up map composed with the projection to the $\mathbb{C}$ factor $y:X\to V\times\mathbb{C}\to\mathbb{C}$ parametrizes a deformation of a smooth generic fiber to a singular fiber containing the exceptional divisor $E$ over $0\in\mathbb{C}$. Recall that the exceptional divisor is the projectivization of the normal bundle to the submanifold blown-up, here $\mathbb{P}(\mathcal{N}_{H\times\{0\}\backslash V\times\mathbb{C}})$ where $\mathcal{N}_{H\times\{0\}\backslash V\times\mathbb{C}}\cong(\mathcal{L}\oplus\mathcal{O}_{\mathbb{C}})|_{H\times\{0\}}\to H\times\{0\}$.
@@
(1.2) $X=\text{Bl}_{H\times\{0\}}V\times\mathbb{C}=\{(x_{1},\ldots,x_{n},y,(u_{1}:u_{2}))\in\mathbb{P}(\mathcal{L}\oplus\mathcal{O}_{\mathbb{C}})\mid s(x_{1},\ldots,x_{n})u_{2}=u_{1}y\}$
@@
Fibers $y^{-1}(t)\cong V$ for $0\neq t\in\mathbb{C}$ because ${x}:=(x_{1},\ldots,x_{n})\in V$ is uniquely determined by $s(x)=t\cdot\frac{u_{1}}{u_{2}}$. As outlined in Remark 1.1, over 0 this fiber deforms to $y^{-1}(0)\cong V\cup_{H}E$ intersecting along $H$ [AAK, Proposition 7.3], as follows. First, the exceptional divisor $E$ is a $\mathbb{P}^{1}$-bundle; letting $p:=\pi|_{E}$ we have
$\text{pr}_{V}\circ p:E\to H.$
Then
(1.3) $E=\pi^{-1}(H\times\{0\})=\{({x},0,(u_{1}:u_{2}))\mid s(x)u_{2}=u_{1}\cdot 0\}\ni({x},0,(u_{1}:u_{2}))\mapsto{x}\in H$
because $s(x)u_{2}=u_{1}\cdot 0$ for all $(u_{1}:u_{2})\in\mathbb{P}^{1}\cong\mathbb{P}(\mathcal{L}_{{x}}\oplus\mathcal{O}_{\{0\}})$ implies $s(x)=0$ therefore $x\in H$. Second, the fiber $y^{-1}(0)$ over 0 also contains the proper transform $\tilde{V}$ of $V\times\{0\}$, which is
$\tilde{V}=\overline{\pi^{-1}((V\backslash H)\times\{0\})}=\{({x},0,(1:0))\mid{x}\in V\}\cong V.$
Intersecting these two sets
@@
To obtain a Lagrangian torus fibration on $X$, start with the moment map on $V\times\mathbb{C}$ obtained from the torus action $(e^{i\theta_{1}}x_{1},\ldots,e^{i\theta_{n}}x_{n},e^{i\theta_{n+1}}y)$. Then use cut and paste operations [4from2] on the moment polytope in the base to blow up $V\times\mathbb{C}$ along $H\times\{0\}$. The resulting polytope is the base of a Lagrangian torus fibration on $X$, with coordinates
(1.4) $(\xi_{1},\ldots,\xi_{n},\eta)=\frac{1}{2\pi}(\log|x_{1}|,\ldots,\log|x_{n}|,\mu_{X}(x_{1},\ldots,x_{n},y))$
for moment map $\mu_{X}$ from the $S^{1}$-action on the $y$ coordinate of the blow-up $X$ [AAK, §4].
Generalized SYZ mirror symmetry. Here is an outline of the generalized SYZ procedure [AAK], which we will later describe in more detail for the setting of this paper. Let
$\Omega_{V\times\mathbb{C}}=i^{n+1}\left(\bigwedge_{j=1}^{n}d\log x_{j}\right)\wedge d\log y$
be a meromorphic $(n+1)$-form on $V\times\mathbb{C}$ with simple poles on the toric divisors $D_{V\times\mathbb{C}}=(V\times 0)\cup(D_{V}\times\mathbb{C})$ where the term $D_{V}$ is not there if $V$ is an abelian variety. Then we have a canonical form on $X$ given by
$\Omega=\pi^{*}\Omega_{V\times\mathbb{C}}$
@@
$D=\tilde{V}\cup\pi^{-1}(D_{V}\times\mathbb{C}).$
@@
The manifold $Y$ can be defined by a polytope if $V$ is toric, or a polyhedron if $V$ is abelian. It is defined in terms of the tropicalization of the defining section $s$ of the hypersurface $H$:
@@
For $V$ an abelian variety there are additional steps to quotient by its defining lattice. Both $X$ and $Y$ admit Kähler structures and the holomorphic functions $y:X\to\mathbb{C}$, $W_{0}:Y^{0}\to\mathbb{C}$, and $v_{0}:Y\to\mathbb{C}$ are Landau-Ginzburg models. In this paper, we take $X$ without superpotential to be the complex side (B-side) of mirror symmetry and $(Y^{0},W_{0})$ to be a symplectic (A-side) Landau-Ginzburg model, see [ACLL_LG, Definition 1.1].
@@ 1.2. Statement of main theorem
Now we assume $V$ is an abelian surface of complex dimension 2 and $H$ is a genus 2 curve of complex dimension 1. Also, the symbol $n$ will now be used to denote an element in $\mathbb{Z}^{2}$, so we let the dimension of $V$ be denoted by $g$. The $g=2$ HMS result of [Ca20] was between $H$ and $(Y,-v_{0})$. Here it is between $X$ and $(Y^{0},W_{0})$.
Let $V=(\mathbb{C}^{*})^{2}/\tau\mathbb{Z}^{2}$ be the principally polarized abelian surface, topologically a 4-torus $T^{4}$, corresponding to the complex modular parameter $\tau=\frac{i}{2\pi}\log t\left(\begin{matrix}2&1\\ 1&2\end{matrix}\right)$, $0<t\in\mathbb{R}_{\ll 1}$, defined by quotienting by
(1.5) $(\tau n)\cdot(x_{1},x_{2})=(e^{2\pi i\sum_{j=1}^{2}\tau_{1j}n_{j}}x_{1},e^{2\pi i\sum_{j=1}^{2}\tau_{2j}n_{j}}x_{2}),\quad(x_{1},x_{2})\in(\mathbb{C}^{*})^{2},\quad(n_{1},n_{2})\in\mathbb{Z}^{2}.$
@@
Define $X:=\text{Bl}_{H\times\{0\}}V\times\mathbb{C}$ with blow-up map $\pi:X\to V\times\mathbb{C}$, exceptional divisor $p:E\to H\times\{0\}$, and inclusion $i:H\times\{0\}\to V\times\mathbb{C}$. To obtain the generalized SYZ mirror, we have an additional step of quotienting by $\tau\mathbb{Z}^{2}$ after restricting $Y$ to a neighborhood where the group acts freely [Ca20, Eqs. (3.10), (3.24), (3.25)]. For small $\varepsilon<1$,
@@
$\Delta_{Y}$ $=(\Delta_{\tilde{Y}})|_{|\chi^{(0,0,1)}|\leq\varepsilon}/\tau\mathbb{Z}^{2}\cong T^{2}\times[0,\varepsilon^{\prime}]$
where $\chi^{(0,0,1)}$ is the complex affine toric coordinate corresponding to the vector in the $\eta$ direction, $(0,0,1)$. We will denote it by $v_{0}$ following [Ca20, AAK]. The polytope $(\Delta_{\tilde{Y}})|_{|\chi^{(0,0,1)}|\leq\varepsilon}$ looks like a bowl tiled by infinitely many hexagons and thickened by an amount $\varepsilon^{\prime}$ that depends on $\varepsilon$. After the quotient, the hexagon is identified to a $T^{2}$ and $\Delta_{Y}\cong T^{2}\times[0,\varepsilon^{\prime}]$.
Recall from toric geometry that each vertex $v$ corresponds to a chart $\text{Spec}(\mathbb{C}[y_{1}^{v},y_{2}^{v},y_{3}^{v}])\cong\mathbb{C}^{3}$ in local inhomogeneous coordinates $y_{1}^{v},y_{2}^{v},y_{3}^{v}$, and edges define transition functions. The reason this polytope defines the SYZ mirror is because its vertices and edges match the SYZ charts and transition functions. In particular,
$v_{0}=y_{1}^{v}y_{2}^{v}y_{3}^{v}.$
Then $(Y,-v_{0})$ is the SYZ mirror to $(X,y)$ and the generalized SYZ mirror to $H$. The SYZ mirror to $X$ is $(Y^{0},W_{0})$ defined by
@@
where $T$ is the Novikov parameter and $\epsilon>0$ is small (different from the $\varepsilon$ in the domain of definition in Equation (1.6)). The Novikov parameter parametrizes complex structures on $Y$ mirroring symplectic structures on $X$, but here we consider complex structures on $X$ mirroring symplectic structures on $Y$, which is parametrized by $\tau$. In general, an SYZ complex mirror is defined over a Novikov parameter $T$ so that structure maps converge in the Fukaya category of the original symplectic manifold. Here, structure maps converge over $\mathbb{C}$ and we do not consider the A-side of the genus 2 curve (which was considered in [genus2]), so we may assume
(1.8) $T=e^{-2\pi}.$
Here $W_{0}:Y^{0}\to\mathbb{C}^{*}$ is a $T^{4}$-fibration over $\mathbb{C}^{*}$ with one singular fiber over $T^{\epsilon}$. We use the symplectic form on $Y$ constructed in [Ca20, §3.5], restricted to $Y^{0}$. The total space of $Y$ is smooth, and $\omega$ is modeled on the standard form for $\mathbb{C}^{3}$ near critical points of $v_{0}$ (the north and south pole of the banana manifold of three $\mathbb{P}^{1}$’s identified at their poles), while it is modeled on the standard product form for $\mathbb{CP}^{2}(3)\times\mathbb{D}$ away from the critical points.
@@
What was called $x,y,z$ in [Ca20, Figure 10, §3.3] is $y_{1}^{(-1,-1,0)},y_{2}^{(-1,-1,0)},y_{3}^{(-1,-1,0)}$ here. And $\tau$ in [Ca20] has been renamed $t$ here, in order to reserve $\tau$ for the more standard notation of the complex modular parameter, here $\frac{i}{2\pi}\log t\left(\begin{matrix}2&1\\ 1&2\end{matrix}\right)$.
To recap, here are the three forms of SYZ mirror symmetry in [AAK] in the case of $V$ an abelian surface.
Theorem 1.4 ([AAK, Theorem 10.4]).
@@
HMS for (3) is examined in [Ca20]. HMS for (2) is the subject of this paper.
@@
[…]
@@
Semi-orthogonal decomposition for blowups. We adapt the diagram at the start of [Or92, §4] where their $\tilde{X}\to X$ is our blow up $\pi:\text{Bl}_{H\times\{0\}}V\times\mathbb{C}\to V\times\mathbb{C}$, their $p:\tilde{Y}\to Y$ is the exceptional divisor which is our $\mathbb{P}^{1}$-bundle $p:E\to H\times\{0\}$, and $j:E\to X$ is inclusion. That is, the following commutes:
@@
Then [Or92, Assertion 4.2(b), Theorem 4.3] assert that there is a semi-orthogonal decomposition
@@
where $\mathcal{O}_{p}(-1)=\mathcal{O}_{E}(E)$. In the first factor, this means that a sheaf on $H$ is pulled back to the exceptional divisor and then extended by 0 to all of $X$.
This result holds, roughly speaking, because of Beilinson’s spectral sequence in [Huy06, Proposition 8.28] and the semi-orthogonality of $\mathcal{O}_{\mathbb{P}^{n}}$ with $\mathcal{O}(1)_{\mathbb{P}^{n}}$. Specifically, $H^{0}(\mathbb{P}^{n},\mathcal{O}(1))\cong\mathbb{C}[x_{0},\ldots,x_{n}]_{1}$, degree 1 homogeneous polynomials in $(n+1)$-variables, but $H^{0}(\mathbb{P}^{n},\mathcal{O}(-1))=0$.
Main Theorem. First we set some notation on the B-side.
Definition 1.6 (notation $D^{b}_{\mathcal{L}}$, [Ca20]).
Let $\text{pr}_{V}:V\times\mathbb{C}\to V$ and $\text{pr}_{\mathbb{C}}:V\times\mathbb{C}\to\mathbb{C}$ be the projections to each factor. We define the following subcategories of $D^{b}\text{Coh}$.
@@
    $D^{b}_{\mathcal{L}}\mathrm{Coh}(X)$: full subcategory of objects in $j_{*}(p^{*}D^{b}_{\mathcal{L}}\mathrm{Coh}(H\times\{0\})\otimes\mathcal{O}_{p}(-1))$ and $\pi^{*}D^{b}_{\mathcal{L}}\mathrm{Coh}(V\times\mathbb{C})$
Now we describe the A-side category. On $(Y^{0},W_{0})$ we take the Fukaya-Seidel category $FS(Y^{0},W_{0})$, after Fukaya [fooo1] for compact manifolds, adapted to exact Lefschetz fibration symplectic LG models in [seidel], and generalized to non-exact, non-Lefschetz LG models in [ACLL_LG]. Then our main theorem is: